### v2

# Refine Neutrino Events Reconstruction with BEiT-3

Abstract
Neutrino Events Reconstruction has always been crucial for IceCube Neutrino Observatory. In the Kaggle competition “IceCube – Neutrinos in Deep Ice”, many solutions use Transformer. We present ISeeCube, a pure Transformer model based on TorchScale (the backbone of BEiT-3). When having relatively same amount of total trainable parameters, our model outperforms the $2^{\textrm{nd}}$ place solution. By using TorchScale, the lines of code drop sharply by about 80% and a lot of new methods can be tested by simply adjusting configs. We compared two fundamental models for predictions on a continuous space, regression and classification, trained with MSE Loss and CE Loss respectively. We also propose a new metric, overlap ratio, to evaluate the performance of the model. Since the model is simple enough, it has the potential to be used for more purposes such as energy reconstruction, and many new methods such as combining it with GraphNeT can be tested more easily. The code and pretrained models are available at https://github.com/ChenLi2049/ISeeCube.
preprint: AIP/123-QED
I Introduction
IceCube is a neutrino observatory that studies the most extreme phenomena in the universe by observing neutrinos Aartsen, Collaboration et al. (2016). Examples include Gamma-ray Bursts, Active Galactic Nuclei, and recently discovered neutrinos emitted by the Milky Way Collaboration*† et al. (2023). Neutrino Events Reconstruction such as direction reconstruction and energy reconstruction are crucial in the observation of neutrinos. Traditionally, to perform the reconstruction, we would use likelihood methods or Machine Learning algorithm that mostly based on Convolutional Neural Network (CNN) Aiello et al. (2020); Abbasi et al. (2021); Eller et al. (2023), or Graph Neural Network (GNN), Abbasi et al. (2022); Søgaard et al. (2022).
In order to encourage the development of innovative solutions to improve the accuracy and efficiency of Neutrino Event Reconstruction, the Kaggle competition “IceCube – Neutrinos in Deep Ice” Eller (2023) was held between January 19, 2023 and April 19, 2023. Most top solutions use Transformer or GraphNeT integrated with Transformer. Transformer is traditionally used for Natural Language Processing Vaswani et al. (2017), and then Computer Vision Dosovitskiy et al. (2020) by introducing the idea of class token (cls token) and only using the encoder part of the full Transformer. Transformer for vision is typically called Vision Transformer (ViT). Transformer is powerful that it can learn the position information of a sentence or a picture.
In particular, the $2^{\textrm{nd}}$ place solution IceMix uses BEiT v2 Peng et al. (2022) block as encoder. BEiT v2 uses semantic-rich visual tokenizer as the reconstruction target for masked prediction in Computer Vision. However, current models are heavy and the benefits of each other is hard to be gathered.
BEiT-3 Wang et al. (2022a) is a general-purpose multimodal foundation model based on Magneto Wang et al. (2022b). And it can be used by importing TorchScale Ma et al. (2022), which is an open-source toolkit that enables scaling Transformers both efficiently and effectively. ISeeCube is a pure Transformer model based on TorchScale. A pure Transformer struture has been proven powerful in many Machine Learning missions, and in Neutrino Events Reconstruction through this competition. And by utilizing the highly integrated TorchScale, it is easier for refining, maintaining and testing.
For upcoming neutrino observatories such as TRIDENT Ye et al. (2023) and HUNT Huang et al. (2023), considering that (a) they also observe astrophysical neutrinos and (b) their structure is close to IceCube that they are composed of several strings of Digital Optical Modules (DOMs), there’s a great chance that ISeeCube can also be used for these observatories.
The paper is structured as follows: Section II summarize Neutrino Events Reconstruction methods based on likelihood principles, CNN and GNN, and some of the top solutions in this Kaggle competition. We present ISeeCube as a pure Transformer structure for Neutrino Events Reconstruction in section III, and compare regression models and classification models. Followed by section IV we describe the dataset and the training process, and we evaluate the azimuthal and zenithal distribution of the prediction made by regression and classification models. In section V we discuss the potential of ISeeCube and challenges in other missions such as energy reconstruction and cascade/track event classification.
II Related Work
In section II.1 we summarize some of the likelihood methods and Machine Learning methods used in neutrino reconstruction, followed by section II.2 we summarize some of the top solutions in the Kaggle competition.
II.1 Previous Neutrino Events Reconstruction Methods
Previously, likelihood method is widely used for Neutrino Events Reconstruction:
  • SkyLLH Bellenghi, Karl, and Wolf (2023) is a Python-based tool that uses log-likelihood functions.
  • Based on Jammy_flows, by using the differential entropy and the KL-divergence to its maximum entropy approximation Glüsenkamp (2023), the results for showers and muons are interpreted.
Also, there are Machine Learning models based on CNNs or GNNs, for example:
  • CNN model combined with hexagonally shaped kernels Abbasi et al. (2021), which takes advantage of the geometric symmetry in the shape of IceCube.
  • A likelihood-free technique Eller et al. (2023) in which only simulated data is required as input without an explicit assumption of the likelihood function.
  • Deep CNN method Aiello et al. (2020) that treat event as an image in the Neutrino Events Reconstruction for KM3NeT/ORCA.
  • GNN method Abbasi et al. (2022) that treat event as point cloud. It is capable of distinguishing neutrino events from cosmic-ray backgrounds, classify different neutrino event types, and reconstructing the deposited energy, direction and interaction vertex.
  • GraphNeT Søgaard et al. (2022), an open-source python package for Neutrino Events Reconstruction based on GNN. One of the advantages of GraphNeT is that it can be implemented on the original model smoothly, since the GNN process can be defined that it does not change the shape of input.
  • A hybrid approach Hünnefeld (2021) that utilizes generative CNN to approximate the likelihood, which may then be used in a traditional maximum-likelihood setting.
  • CNN model Peterson, Rodriguez, and Hanson (2023) that exploits time and depth translational symmetry in IceCube DeepCore that distinguishing muon neutrinos from other flavors.
Neutrinos from the Milky Way is detected Collaboration*† et al. (2023) by first using CNN Abbasi et al. (2021) to classify cascade events from all the events, and then using hybrid reconstruction method combined with a maximum likelihood estimation. The hybrid reconstruction method uses Neural Network (NN) including symmetries and detector-specific domain knowledge, which is described previously on Hünnefeld (2021), and further discussed in the supplementary materials of Collaboration*† et al. (2023).
II.2 In the Kaggle Competition
In private leaderboard of the Kaggle competition  [^1] https://www.kaggle.com/competitions/icecube-neutrinos-in-deep-ice/leaderboard, these solutions are representative and quite interesting.
  • $1^{\textrm{st}}$: It’s an combination of EdgeConv and Transformer. The Transformer part only has 3-4 layers. The number of total parameters is only 6M, which is the smallest top solutions.
  • $2^{\textrm{nd}}$: The Transformer part is based on BEiT v2 Peng et al. (2022) block. Also, it uses class token and introduces relative spacetime interval bias as Relative Positional Embedding (RPE) Shaw, Uszkoreit, and Vaswani (2018). This solution shows that, with right set of parameters for RPE, embedding dimensions and number of layers, a pure transformer structure works as well. RPE is close to the mechanism that is used in GNN to capture the relationship of nodes in a sequence.
  • $3^{\textrm{rd}}$: It’s a pure Transformer structure, which has two classification heads (for azimuth and zenith) with 128 bins each. Since it’s a classification mission, a Cross Entropy Loss is better than Von Mises-Fisher Loss Kumar and Tsvetkov (2018).
  • $5^{\textrm{th}}$: The GraphNeT part uses LayerNorm as the activation function instead of BatchNorm1d. The Transformer part has two classification heads, each containing $256$ bins for azimuth and zenith. The Transformer consists of $8$ layers, with a model size of $512$ and $8$ attention heads. Additionally, a $3$-layer Bi-directional LSTM model selector is employed to choose between GraphNeT and Transformer. Combining multiple individual models to solve a given problem is called “ensemble” method.
Besides GNN, which wins the Early Sharing Prize and is provided as an example in this competition, some solutions also use Recurrent Neural Network (RNN), which takes into account of time information. Long Short-Term Memory (LSTM) is designed to deal with long-range dependencies. Bidirectional Gate Recurrent Unit (BGRU) is designed to deal with the same problem, with better performance. But as Vaswani et al. (2017) stated, the attention mechanism can be used to solve the same problem and reach better score.
Table 1: Different Machine Learning methods used in some of the Kaggle solutions.
Transformer GNN RNN Loss Ensemble
$1^{\textrm{st}}$checkmark EdgeConv VMF (custom) $\checkmark$
$2^{\textrm{nd}}$ BEiT v2 EdgeConv (optional) VMF $\checkmark$
$3^{\textrm{rd}}$ NanoGPT CE $\checkmark$
$5^{\textrm{th}}$checkmark GraphNeT LSTM CE $\checkmark$
$6^{\textrm{th}}$checkmark $\checkmark$ LSTM CE (custom) $\checkmark$
$8^{\textrm{th}}$ GraphNeT, GPS, GravNet VMF, CosineSimilarity
$10^{\textrm{th}}$ LSTM CE, VMF
$11^{\textrm{th}}$checkmark $\checkmark$ GRU $\checkmark$
$12^{\textrm{th}}$ DynEdge LSTM $\checkmark$
$13^{\textrm{th}}$ GraphNeT $\checkmark$
$14^{\textrm{th}}$checkmark BGRU
$15^{\textrm{th}}$ DynEdge VMF
$17^{\textrm{th}}$ DynEdge LSTM $\checkmark$
$20^{\textrm{th}}$ LSTM $\checkmark$
Table 1 summarizes some of the top solutions in the Kaggle competition Bukhari et al. (2023). $\checkmark$” means the architecture is used in the solution, and a specific name represents the model of the architecture. “Ensemble” means whether the final result is a ensemble of several different models by some weighed mechanism.
In summary, Transformer architecture can work really well for having the ability to capture the relationship of both nodes Joshi (2020) and time sequence. What’s more, since the total number of nodes is smaller than vocabulary of a language, ViT is better than Transformer that an event is more close to a picture than sentence.
As for loss function, Von Mises-Fisher Loss Kumar and Tsvetkov (2018) has the advantage that the first Bessel function is the solution for partial differential equations in a cylinder. And IceCube is constructed by several strings, meaning that all the detectors on the same string has the same $x$ and $y$ coordinate, which makes its shape more close to a cylinder. This loss function finds the internal properties of IceCube. However, some solutions indicate that Cross Entropy Loss is better than Von Mises-Fisher Loss. A possible explanation is that these solutions divide the entire sky into several bins and then predict which bin is the most likely, thus it can work really well with Cross Entropy Loss, which is normally used for classification.
III ISeeCube
We present a pure Transformer model for direction reconstruction: ISeeCube. ISeeCube has the potential to be extended to other tasks, such as energy reconstruction, cascade/track event classification, etc. In this section, we will elaborate the structure of ISeeCube:
(a) ISeeCube
(b) Transformer Encoder
Figure 1: The structure of ISeeCube. (a) shows how raw data is embedded then put into the encoder, the class token is taken out and transformed into azimuth and zenith as the result, while the register tokens and embedded data are removed. “n” in “emb n” corresponds to the number of pulses in an event, while “n” in “reg n” is a rather arbitrary number. (b) shows the detail of each layer in the Transformer encoder, which is originally figure 2 (a) of Wang et al. (2022b).
  • In section III.1, we summarize previous data representation of neutrino events used in CNNs and GNNs, and conclude that Embedding used in Transformer is a intuitive derivation of the previous methods.
  • In section III.2 we show that by using TorchScale, the lines of code drop sharply. We show that ViT-like model is capable of learning the position and time information of neutrino events, and future new models or upgrades on Transformer Encoder and ViT has the potential to be used on Neutrino Events Reconstruction. Simple methods work really well, such as Sinusoidal Position Embedding, Relative Position Embedding (RPE) Raffel et al. (2020), Class Token and Register Tokens Darcet et al. (2023).
  • In section III.3 we compare regression and classification models by using different Multilayer Perceptron (MLP) on the class token.
  • In section III.4 we compare Mean Squared Error (MSE) Loss and Cross Entropy (CE) Loss used in regression and classification models, respectively. We also discuss the usage of the official metric in the Kaggle competition.
III.1 Embedding: Data Representation
Machine Learning models in Computer Vision mostly focus on images that satisfy the following properties:
  • Image $\mathbf{x}$ has the shape $\mathbb{R}^{H\times W\times C}$. This means that (a) the image is 2-dimensional, thus $H$ (Height) and $W$ (Width), and (b) the image has a grey channel ($C=1$), or RGB channels ($C=3$).
  • Image is compactly arranged, meaning that the distance between each pixels is the same. And in most cases, zero.
On the other hand, in IceCube, the neutrino events satisfy the following properties:
  • Intuitively, considering all the DOMs (no matter detect photons in this particular event or not), neutrino event can be represented as a 3-dimensional “image” $\mathbb{R}^{H\times W\times L\times P}$. This means that (a) neutrino event is 3-dimensional, thus $H$ (Height), $W$ (Width) and $L$ (Length), and (b) each pulse (detection made by DOM) has several properties, in our case time, charge and auxiliary ($P=3$).
  • Neutrino event is loosely arranged, meaning that (a) the distance between each DOM is not the same, and (b) not all the DOMs detect photons in an event, see figure 2
One of the problems of this intuitive representation might be computation because (a) this matrix is much larger than images in traditional Machine Learning methods and (b) large part of the matrix is empty (sparse matrix). Thus in many previous studies the intuitive representation is not used. In the rest of this section we will review some of the methods used in CNNs, GNNs and Transformers.
CNNs: In Peterson, Rodriguez, and Hanson (2023) and Abratenko et al. (2021), neutrino event is represented as $\mathbb{R}^{S\times N\times T}$, where $S$ is the number of strings, $N$ is the number of DOMs on each string and $T$ is time segments. Time segments is the total time divided by time periods. And the number in this matrix represents charge. This matrix is then put into a CNN. For example, in Abratenko et al. (2021) the CNN is based on ResNet He et al. (2016).
GNNs: In GraphNeT and many other GNN-based methods, the neutrino event is represented as $\mathbb{R}^{N\times P}$, where $N$ is the number of DOMs that detect photons in a particular event, $P$ is the number of properties a pulse has, in our cases position ($x$, $y$, $z$), time, charge and auxiliary ($P=6$). This matrix is then put into a GNN. For example, in GraphNeT, TAGConv and EdgeConv from the package PyTorch Geometric are used.
Transformers: Same as GNN method, the neutrino event is also represented as $\mathbb{R}^{N\times P}$. Firstly, when the total number of pulses is bigger than the given $N$, we would choose auxiliary False first, then True; when the total number of pulses is smaller than $N$, the empty space will be filled with the position of the first DOM (the choice of the first DOM is rather arbitrary) and time, charge, auxiliary would be zero. For more information on the dataset and how to normalize the event, see section IV.1. Then Sinusoidal Position Embedding is used. In Vaswani et al. (2017), Sinusoidal Position Embedding assigns a unique fixed vector to each position in the input sequence. These vectors are computed using sine and cosine functions with different frequencies. In neutrino events it is introduced for the same purpose, projecting $\mathbb{R}^{N\times P}$ to a higher dimension. In our case, for example, one of the properties $x$ (shape $\mathbb{R}^{N\times 1}$), is projected to $\mathbb{R}^{N\times N}$ by:
$PE_{(i,j)}$ $=\sin\left(\frac{pos(i)}{10000^{j/N}}\right),j=0,1,\cdots,\frac{N}{2}-1$ (1)
$PE_{(i,j)}$ $=\cos\left(\frac{pos(i)}{10000^{j/N}}\right),j=\frac{N}{2},\frac{N}{2}+1,\cdots,N-1$
where $pos(i)$ is each element in $\mathbb{R}^{N\times 1}$. Then the rest of the properties will go through the same process and all the embedded elements will be concatenated together ($x$, $y$, $z$ will be concatenated first, then the rest of the properties). Then we use a MLP to produce the final embedded event as $\mathbb{R}^{N\times d_{model}}$. This is the input of the Transformer Encoder.
As we can see, Embedding used in Transformers is an intuitive upgrade of GNNs. Also, normally, when using higher embedding dimension, we will get better performance, in which way we can weigh the balance between computation and performance more easily. While in GNNs, different structures have to be tested multiple times before we can find the most suitable structure for specific model and goal.
III.2 Transformer Encoder
We present an easy-to-use and simple backbone structure with lines of code dropping sharply by about 80%. TorchScale is portable and easy to maintain:
  • Out-of-the-box: An encoder structure can be used by few lines of code. A lot of methods can be tested by adjusting the config. For example, Relative Position Embedding (RPE). While in the $2^{\textrm{nd}}$ place solution, the layers for encoder have be rewritten.
  • Only relay on few dependencies including PyTorch, timm and fairscale, while GNNs rely on PyTorch, PyTorch Cluster, PyTorch Scatter, PyTorch Sparse and
    PyTorch Geometric.
  • In Feed Forward Network (FFN), A sub_LayerNorm (sub_LN) is added to stabilize models with deeper layers. Also, the default activation function in FFN is now GELUs Hendrycks and Gimpel (2016) rather than ReLUs, which combines dropout zoneout and ReLUs to alleviate the vanishing gradient problem. Also, a dropout layer is added to alleviate the same problem.
Relative Position Embedding (RPE) Raffel et al. (2020) is presented as a new method to learn the relative position information. In the $2^{\textrm{nd}}$ place solution, RPE is added by introducing Special Relativity bias and writing new layers and blocks with this bias. In TorchScale, RPE can be added easily by adjusting few parameters, though without Special Relativity bias.
In ViT Dosovitskiy et al. (2020), Class Token is concatenated to the embedding of the picture to learn the position information. In ISeeCube, Class Token is concatenated to the embedding of a neutrino event for the same purpose, to learn the position information of neutrino events.
Register Tokens Darcet et al. (2023) are used for smoother feature maps and attention maps for downstream visual processing when finetuning ViT models. In ISeeCube, Register Tokens are concatenated to the embedding of a neutrino event for the same purpose. Also, we notice that when the number of class token is $1$, the number of register tokens is $3$, and we pick $196$ pulses in an event, the total sequence length been put into the encoder is $200$. The “warp size” of a CUDA core is the number of threads in a warp executed simultaneously on a CUDA core, which is normally $32$. When multiple threads within a warp access adjacent memory locations, the GPU can coalesce these memory transactions into a single memory transaction, resulting in improved memory throughput. $200$ is a multiple of $8$, which is more CUDA-friendly than the case with only class token, $197$, which is a prime number.
III.3 Multilayer Perceptron: Regression and Classification
For a continuous space, regression and classification are commonly used methods to predict the output of a neural network. We use different Multilayer Perceptron (MLP) as the final layer to build these two models:
  • Regression: The model predicts the output itself. The output is produced by a MLP that results with a tensor with shape $[B,3]$, where $B$ is mini-batch size and 3 represents $x_{pred}$, $y_{pred}$, $z_{pred}$, which belong to $[-1,1]$.
  • Classification: We classify the continuous space into several bins then the model predicts which bins does the prediction belongs to. The predicted azimuth is produced by a MLP that results with a tensor with shape $[B,N_{bins}]$, where for azimuth, we use a MLP head with $N_{bins}=128$; for zenith, we use another MLP head with $N_{bins}=64$, so that the width for both bins is $\frac{\pi}{64}$. Then we use softmax to choose the largest index, which corresponds to an angle.
We trained three models (The hyper-parameter is summarized in table 3. For detailed training configuration and loss curve, please see section IV.2.):
  • Model S-RegA is a regression model using the structure of figure 1.
  • Model S-RegB is a regression model using similar structure, replacing one encoder with two encoders with and without Relative Position Bias, and residual connections are used bewteen these two encoders.
  • Model S-ClsB is a classification model using the same structure as model S-RegB.
We tried to train a classification model using the same structure as model S-RegA, but the loss does not go down. This is one of the disadvantages of classification model that they tend to be more unstable.
III.4 Loss
We trained the regression model using Mean Squared Error (MSE) Loss, and the classification model using Cross Entropy (CE) Loss. The competition Metric is used as metric in the validation process.
III.4.1 Mean Squared Error Loss
As stated in section II.2, the shape of IceCube is close to a cylinder, and the first Bessel function is the solution for partial differential equations in a cylinder, thus the Von Mises-Fisher Loss works really well. But in ISeeCube, a more primitive loss function such as Mean Squared Error (MSE) Loss can work just well.
Mean Squared Error (MSE) Loss is a common loss function for regression models. It calculates mean squared error between each element in the prediction $x$ and target $y$:
$\textrm{MSE}\left(x,y\right)=\frac{1}{n}\sum_{i}\left(x_{i}-y_{i}\right)^{2}$ (2)
where $x_{i}$ and $y_{i}$ can be in any shape. In our case, $x_{i}$ and $y_{i}$ are both one-dimensional array with length $3$ (representing $x_{\text{reco}}$, $y_{\text{reco}}$, $z_{\text{reco}}$ and $x_{\text{true}}$, $y_{\text{true}}$, $z_{\text{true}}$ respectively).
Since MSE Loss is a rather primitive loss function, it could be used for more tasks such as energy reconstruction. And to prevent exploding gradients, it can be replaced by Smooth L1 Loss Girshick (2015).
III.4.2 Cross Entropy Loss
And for classification models, following the Kaggle solutions stated in II.2, we use another rather primitive loss function Cross Entropy (CE) Loss:
$\textrm{CE}\left(x,y\right)=\frac{1}{n}\sum_{i}x_{i}\log{y_{i}}$ (3)
where $x_{i}$ is every item in the predicted probability and $y_{i}$ is the true probability index. We perform CE Loss on two directions, azimuth and zenith, respectively. And for each angle, $x$ and $y$ are one-dimensional array with length $N_{bins}$.
Similarly, since CE Loss is a rather primitive loss function, it could be used for more tasks such as cascade/track classification.
III.4.3 Competition Metric
The derivation of Competition Metric Eller (2023) (Mean Angular Error) is as follows. In spherical coordinate system, given two vectors true direction $\vec{r}_{\textrm{true}}$ and predicted direction $\vec{r}_{\textrm{reco}}$ with radius equals $1$ (after normalization), we have:
$\vec{r}_{\text{true}}$ $=\hat{r}+\hat{\theta}\vartheta_{\text{true}}+\hat{\phi}\varphi_{\text{true}}$ (4)
$\vec{r}_{\text{reco}}$ $=\hat{r}+\hat{\theta}\vartheta_{\text{reco}}+\hat{\phi}\varphi_{\text{reco}}$
In 3D Cartesian coordinate system, after normalization (with $r_{\textrm{reco}}=1$):
$x_{\text{reco}}$ $=\sin{\vartheta_{\text{reco}}}\cos{\varphi_{\text{reco}}}$ (5)
$y_{\text{reco}}$ $=\sin{{\vartheta_{\text{reco}}}}\sin{{\varphi_{\text{reco}}}}$
$z_{\text{reco}}$ $=\cos{\vartheta_{\text{reco}}}$
these two vectors are transformed into:
$\vec{r}_{\text{true}}$ $=\hat{i}\sin{\vartheta_{\text{true}}}\cos{\varphi_{\text{true}}}+\hat{j}\sin{{\vartheta_{\text{ture}}}}\sin{{\varphi_{\text{true}}}}+\hat{k}\cos{\vartheta_{\text{true}}}$ (6)
$\vec{r}_{\text{reco}}$ $=\hat{i}\sin{\vartheta_{\text{reco}}}\cos{\varphi_{\text{reco}}}+\hat{j}\sin{{\vartheta_{\text{reco}}}}\sin{{\varphi_{\text{reco}}}}+\hat{k}\cos{\vartheta_{\text{reco}}}$
Thus the angular error is the arccosine of the inner product of these two vectors:
$\Psi=\arccos(\sin{\vartheta_{\text{true}}}\sin{\vartheta_{\text{reco}}}(\cos{{\varphi_{\text{ture}}}}\cos{{\varphi_{\text{reco}}}}+\sin{{\varphi_{\text{ture}}}}\sin{{\varphi_{\text{reco}}}})+\cos{\vartheta_{\text{true}}}\cos{\vartheta_{\text{reco}}})$ (7)
And Mean Angular Error on the entire test dataset is the average of $\Psi$ for every event in the entire dataset. Random guessing will get a result of $\frac{\pi}{2}$, when two vectors are orthogonal.
IV Experiments
In section IV.1 we summarize some of the most important aspects of the dataset and the normalization of neutrino events. In section IV.2 we show the learning rate schedule, loss curve and performance of our models, followed by section IV.3 we further explore the performance of our models by comparing azimuthal and zenithal distribution. Also in this section we propose a new metric: Overlap Ratio.
IV.1 Dataset
Each event in the dataset is composed of a series of pulses. Most neutrino events have less or equal to $200$ pulses, and in most Kaggle solutions the dimension of the model is less than $200$, meaning that if the number of pulses is bigger than the dimension, some of the pulses will be discarded.
Figure 2: In batch one, 94.3% of events have less or equal to 200 pulses. Most events have about $50$ pulses. This figure is originally used in https://www.kaggle.com/code/rasmusrse/graphnet-example.
[…]
Figure 3: This is an example event where auxiliary False and True are shown separately. It is intuitive that False will get better result in terms of shape thus direction reconstruction. This figure is originally used in Eller (2023).
The biggest challenge of the dataset is its bigness (more than $138$ million events distributed in $660$ batches with the size of over 100 GBs), so we use the same chunk-based training method used in the $2^{\textrm{nd}}$ place solution, meaning that every epoch is trained on $654$ batches divided into $8$ sub-epochs, see section IV.2. The advantage of this chunk-based method is that now we can train the model on the entire dataset with relatively low memory cost. However it may have a disadvantage that a random sampling can only sampling across the sub-set of the entire dataset, which may leads to bias.
IV.2 Training and Performance
For comparison, the hyper-parameters for $2^{\textrm{nd}}$ place solution and ISeeCube are shown in table 2 and table 3. In order to better compare the performance of different Embedding methods and Transformer Encoders, the models that combines Transformers and GNNs in the $2^{\textrm{nd}}$ place solution is not included in table 2.
Table 2: Hyper-parameters and score of $2^{\textrm{nd}}$ place solution without the models that are a combination of Transformer and GNN. “T”, “S” and “B” refers to “Tiny”, “Small” and “Base”. “d32” or “d64” means the dimension of the head is 32 or 64. CV512 means the number of pulses in an event is 512 when inferencing on the validaiton dataset (batch 655 $\sim$ batch 659). Public LB and Private LB refers to the score in the Kaggle competition using public test dataset and private test dataset. The score for competition metric is the smaller the better.
model $d_{model}$ heads layers params CV512 Public LB Private LB
T d32 192 3/6 4+12 7.57M 0.9704 0.9693 0.9698
S d32 384 6/12 4+12 29.3M 0.9671 0.9654 0.9659
B d32 768 12/24 4+12 115.6M 0.9642 0.9623 0.9632
B d64 768 12/24 4+12 115.6M 0.9645 0.9635 0.9629
Table 3: Hyper-parameters and score of ISeeCube. “S-RegA”, “S-RegB” and “S-ClsB” refers to different “Small” models (see section III.3). The score for competition metric is the smaller the better.
model $N$ $d_{model}$ $d_{ffn}$ heads layers rel pos max params score
buckets rel pos
S-RegA 196 384 1536 12 16 32 256 31.7M 0.9651 (CV256)
S-RegB 196 384 1536/1536 12/12 6+12 32/0 196/0 35.1M 0.9757 (CV196)
S-ClsB 196 384 1536/1536 12/12 6+12 32/0 196/0 43.6M 1.0095 (CV196)
Model S-RegA follows the learning rate schedule as table 4. Model S-RegB and S-ClsB follows the learning rate schedule as table 5. The reason for epoch 0 in table 4 is that we first train our model for 1 epoch to test the performance and we think load this pre-trained model for further training won’t lead to overfitting, considering that (a) the score is relatively high and (b) there are following 7 epochs with different dataset (see chunk-based method mentioned in section IV.1). In order to save time and resources when training, for epoch 0$\sim$48, we use $N=196$ when loading the data (the pulses in an event is 196 after normalization), and for epoch 49$\sim$56, we use $N=256$. In table 4 and table 5, “COMB” loss means combining two losses by adding them up:
$COMB$ $=\Psi+0.05\cdot MSE$ (8)
$COMB$ $=\Psi+0.05\cdot CE$
where $\Psi$ is competition metric in the Kaggle competition. “div” and “div_final” are hyper-parameter used in Leaner.fit_one_cycle with 1cycle police Smith and Topin (2019) in fastai.
Table 4: Learning rate schedule of model S-RegA.
epoch 0 1$\sim$8 9$\sim$16 17$\sim$24 25$\sim$32 33$\sim$40
loss MSE MSE MSE MSE COMB COMB
lr_max 1e-4 1e-4 1e-4 1e-5 1e-5 0.2e-6
div, div_final 25 25 25 25 25 default
Table 5: Learning rate schedule of model S-RegB and S-ClsB.
epoch 0$\sim$7 8$\sim$15 16$\sim$23 24$\sim$31 32$\sim$39 40$\sim$47
loss MSE/CE MSE/CE MSE/CE COMB COMB COMB
lr_max 1e-5 1e-5 1e-5 1e-5 0.5e-5 0.35e-5
div, div_final 25 25 25 25 15 10
The loss curve of S-RegA, S-RegB and S-ClsB are shown in figure 4. As we can see, the training loss curve is flucuating while the validation loss curve is constantly reducing. Also, by replacing MSE loss with COMB loss, the loss curve drops in a single epoch.
(a) Previous training epochs with MSE/CE loss.
(b) Latter training epochs with COMB loss.
(c) Entire validation epochs with competition metric.
Figure 4: The loss curve of ISeeCube. (a) shows the previous training epochs using MSE/CE loss; (b) shows the latter training epochs using COMB loss; (c) shows the entire validation epochs with competition metric.
As shown in table 2 and table 3, by adding merely 2.4M parameters, our best model S-RegA outperforms $2^{\textrm{nd}}$ place solution’s S d32 model and approaches its B d32 model. Also, compared with CV512, our model uses CV256, which means that our model requires half the GPU memory when inferencing.
Figure 5: By using voting rule as the ensemble method, our model reaches a score of 0.9645.
By using voting rule as the ensemble method we reaches a score of 0.9645, which would be $3^{\textrm{rd}}$ place solution in the Kaggle competition. Though as shown in table 2, the score for Public LB and Private LB tend to be better than the score of CV. Also, by traing a B model with roughly 120M parameters and using ensemble, our model could reach state-of-the-art.
IV.3 Azimuthal and Zenithal Distribution
(a) Azimuth ($\varphi$) distribution.
(b) Zenith ($\theta$) distribution.
(c) Azimuthal and zenithal error of ensemble ($\Delta\varphi$ and $\Delta\theta$).
(d) Azimuthal and zenithal error of ensemble ($\Delta\varphi$ and $\Delta\theta$), zoomed in.
Figure 6: Inference on batch 655. (a) shows the distribution of azimuth; (b) shows the distribution of zenith. The percentage in (a) and (b) refers to overlap ratio between prediction and true value. (c) shows the distribution of azimuthal and zenithal error of the ensemble, while (d) is zoom in of (c).
Figure 6 shows the azimuthal and zenithal distribution of the prediction. For azimuthal distribution, we notice an interesting phenomenon that the prediction has 6 peaks corresponding to the shape of IceCube, which means that our model is effected by the nuance of the geometric shape of IceCube. We define overlap ratio:
$\rho=\frac{\text{overlap area of the prediction and the true value}}{\text{area of the true value}}$ (9)
As shown in figure 6, the overlap ratio of both regression model and classification model are over 90%.
On the contrary, the zenithal distribution is irregular and has lower overlap ratio. The predictions made by regression models tend to be drifted towards $\theta=0$, while the prediction made by classification tends to be drifted towards $\theta=\frac{\pi}{3}$ and $\theta=\frac{\pi}{2}$. Regression models tend to have a higher overlap ratio than classification models.
As we can see from table 3 and figure 6, this ratio uniformly reflects the performance of the model. By fitting the score of the competition metric and the overlap ratio with four data points (S-RegA, S-RegB, S-ClsB and ensemble), we can get figure 7. The curve for zenith ($\theta$) is more linear than the curve for azimuth ($\varphi$), this is most likely due to the fact that azimuth distribution is more uniform than zenith distribution, as shown in figure 6.
We use some of the most common loss in machine learning to investigate the fundamental models for Neutrino Events Reconstruction. The bias of two major models (Regression and Classification) is like two major interpretations of black-body radiation (Wien’s law and Rayleigh–Jeans law), meaning that each prediction has its own bias and we are still looking for a unified model. By using ensemble, we can lower the competition metric score, but the distribution of zenith is largely close to regression. Thus regression is not the unified model.
Figure 7: Fitting competiton metric score with overlap ratio. We can see that overlap ratio uniformly goes up with better competition metric score.
V Discussion and Conclusion
Neutrino Events Reconstruction has always been crucial for neutrino observatories such as IceCube and KM3NeT/ORCA, which traditionally uses CNN or GNN. In the Kaggle competition, many solutions use Transformer. And it is shown that a pure Transformer model performs really well for Neutrino Events Reconstruction.
Compared with GNN integrated with Transformer or an ensemble of several different models including Transformer, GNN and RNN, a pure Transformer structure is easier for refining, maintaining and testing. And integrating other models with this model would be much easier. In this spirit, we build ISeeCube based on the $2^{\textrm{nd}}$ place solution, by replacing BEiT v2 with TorchScale (the backbone of BEiT-3). We also uses Register Tokens, improving the performance while only slightly increase the computation cost. As a result, the lines of code drop sharply and a lot of new methods could be tested by simply adjusting the config or using other models as the first layer. When having relatively same amount of total parameters, our model outperforms the $2^{\textrm{nd}}$ place solution.
We use rather primitive loss functions such as MSE loss and CE loss in order to explore the fundamental models for Neutrino Events Reconstruction. We then investigated how azimuthal and zenithal distribution would be effected by the choice of regression model or classification model. We show that Transformer, especially Vision Transformer has the potential to learn the information of nodes and time series in neutrino events. Future upgrades on Transformer Encoder or Vision Transformer can be used for Neutrino Events Reconstruction.